• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name B. Quinn

    Log entry time 16:27:25 on May14,2010

    Entry number 322532

    keyword=Day shift summary

    No beam all day, restricted access to work on target vacuum and on spectrometer magnet power supplies.

    15:25 J.P. called to say he is beginning to cool target. There may be target alarms.